# Topological Kleene Field Theories: A new model of computation

### v1

@@ 4. Categorical interpretation: Topological Kleene Field Theories
  • $2$-Morphisms: Given two morphisms $f:\mathbb{N}^{\sqcup m}\dashrightarrow\mathbb{N}^{\sqcup m^{\prime}}$ and $g:\mathbb{N}^{\sqcup m}\dashrightarrow\mathbb{N}^{\sqcup m^{\prime}}$, a unique $2$-morphism $f\Rightarrow g$ between them exists if $f$ extends $g$, i.e., if $D_{g}\subseteq D_{f}$ and $f|_{D_{g}}=g$.
@@
In the same spirit, fixed $n\geq 2$, we can consider the category $\textup{Bord}_{n}^{\textup{dy}}$ of dynamical $n$-bordisms as the category given as follows:
@@
  • Morphism: A morphism $(M_{1},X_{1},\bm{\iota}_{1})\to(M_{2},\bm{\iota}_{2})$ between marked dynamical germs is an equivalence class of dynamical bordisms $(W,X)$ between them. Two such bordisms $(W,X)$ and $(W^{\prime},X^{\prime})$ are declared to be equivalent if there exists a boundary preserving diffeomorphism $F:W\to W^{\prime}$ such that $F_{*}(X)=X^{\prime}$.
@@
$Z=Z_{0}|_{\mathcal{B}}:\mathcal{B}\longrightarrow\textup{PRF}.$
@@
is surjective or, in other words, any partial recursive function is representable through a dynamical bordism. Finally, the condition of being surjective on objects reduces to the fact that $\mathcal{B}$ contains the connected objects.
In this context, Theorems 3.9 and 3.15 can be rephrased in this categorical language as follows.
@@
Topological Kleene Field Theories exist. To be precise, if $\textup{Bord}_{n}^{\textup{c-dy}}$ is the wide subcategory of $\textup{Bord}_{n}^{\textup{dy}}$ of clean dynamical bordisms, then
$Z=Z_{0}|_{\textup{Bord}_{n}^{\textup{c-dy}}}:\textup{Bord}_{n}^{\textup{c-dy}}\longrightarrow\textup{PRF}$
is a Topological Kleene Field Theory.
@@
By Theorem 3.15, the reaching function of any clean dynamical bordism is computable, so $Z_{0}$ restricts to a function with values in PRF on $\textup{Bord}_{n}^{\textup{c-dy}}$. Notice that this functor $Z=Z_{0}|_{\textup{Bord}_{n}^{\textup{c-dy}}}$ is surjective on objects since $\textup{Bord}_{n}^{\textup{c-dy}}$ contains the connected objects. Additionally, by Theorem 3.9, given a partial recursive function, there exists a clean dynamical bordism of the standard disc representing it, which implies that $Z$ is full. ∎
@@ 5. Conclusions
The work presented in this paper establishes a new connection between dynamical systems and computability by representing computable functions through the flows of vector fields on bordisms. These ideas raise interesting questions to be explored in future research. First of all, as a result of this work, we have proven that TKFTs represent a fully functional model of computation, on the same footing as Turing machines and partial recursive functions. However, the dynamical nature of the construction enables new perspective to address classical problems of computation through a novel glass.
In this direction, in [10, Corollary 3.3], it was shown that through area-preserving diffeomorphisms of the disc it is possible to encode a larger class of Turing machines, known as Turing machines with (polynomial) advice. These are machines that are immediately provided at the beginning of the computation with an auxiliary string that depends only on the length of the input, and works as an ‘advice’ that allows the machine to transcend the classical limits of computation. Indeed, the advice itself may not be computable, leading to Turing machines able to process some non-computable functions. But, perhaps more interesting, we can also restrict ourselves to computable advices, which can be used to speed-up the computation since the advice is immediately provided. In that case, we expect that these diffeomorphisms of the disc encoding a computable advice can be used as new pieces of basic bordisms that would improve the computational efficiency of clean dynamical bordisms, due to the presence of advice. This emerging feature of clean dynamical bordism would demonstrate that TKFT already exceeds the efficiency of Turing machines and, and is at least as powerful as Turing machines with polynomial advice.
Furthermore, the method developed in this paper links the complexity of the computable function with the topology of the bordism representing it. Indeed, the method of proof creates a bordism that has as strong deformation retract the graph underlying a Turing machine computing the function, so the homology invariants of the bordisms, and in particular its first Betti number, are forced to agree with the ones of the graph. In this direction, a question arises: is it possible to (algorithmically) construct a clean dynamical bordism computing a partial recursive function $f$ with strictly smaller first Betti number than that of any Turing machine computing $f$? This is particularly interesting taking into account that TKFTs are equivalent to Turing machines, since using clean bordisms we will be able to compute exactly the same partial recursive functions, but potentially improving the complexity of the calculation.

### v2

@@ 3.3. Partial functions from dynamical bordisms
4. Proof of the main theorem
In this section, we shall prove Theorem 3.9 by showing that any finite state machine, as described in Section 2.3, can be represented through a dynamical bordism. For simplicity, we shall focus on the case $n=2$ of surfaces, but the constructions can be straightforwardly extended to the higher dimensional case.
4.1. Reversible Turing machines
$\Delta_{M}:\mathcal{S}\to\mathcal{S}$
Theorem 4.1.
Proposition 4.2.
For every multi-tape reversible Turing machine, there exists a single-tape reversible Turing machine computing the same function.
Proof.
4.2. The basic bordism
Before proceeding with the proof of the main result, let us show that we can represent the basic read-write-shift operation as a bordism using elementary operations. For simplicity, throughout this section we will consider the case in which we have only two symbols in our alphabet, 0 and $1$, but it is straightforward to adapt it to work with more symbols.
$s_{\epsilon}:\Lambda\to\Lambda,$
given by $s_{\epsilon}(t)]_{n}=t_{n+\epsilon}$ for all $n\in\mathbb{Z}$. In other words, for $\epsilon=1$, $s_{1}$ shifts the sequence to the left (or the origin to the right); whereas for $\epsilon=-1$, $s_{-1}$ shifts the sequence to the right (or the origin to the left). It is well known that $s_{\epsilon}$ can be extended to an area-preserving diffeomorphism of the disk $\mathbb{D}$, which is equivalent to the celebrated horseshoe map (recall that $\mathbb{D}$ denotes the standard disk of radius $2$). We include a proof for the sake of completeness.
Lemma 4.3.
$\varphi_{\epsilon}:\mathbb{D}\to\mathbb{D}$
Proof.
Let $I_{\varepsilon}=[-\varepsilon,1+\varepsilon]^{2}$ for $\varepsilon>0$ small. Divide the square $I_{\varepsilon}$ into four regions, $X$, $Y$, $Z$ and $T$, as shown in Figure 7. The points of $\Lambda$ inside each region correspond to those two-sided sequences of the form
[…]
Here, the subscript “0” denotes the position of the 0-th digit in the two-sided sequence. Now, through a linear transformation, we map these regions into the new regions $X^{\prime}$, $Y^{\prime}$, $Z^{\prime}$ and $T^{\prime}$, as displayed in Figure 7. Notice that the points of the Cantor set in these regions correspond to sequences of the form
[…]
Since the new regions are pairwise disjoint, it is easy to extend the aforementioned linear transformation to an area-preserving diffeomorphism of the whole disk $\varphi_{+}:\mathbb{D}\to\mathbb{D}$ so that it is the identity on the boundary, and $\varphi_{+}|_{\Lambda}=s_{+}$ (see [13, Proposition 5.1] for details). Taking $\varphi_{-}=\varphi_{+}^{-1}$, we get and area-preserving diffeomorphism with $\varphi_{-}|_{\Lambda}=s_{-}$.
Now, observe that any diffeomorphism of the disk that is the identity on the boundary is isotopic to the identity with a boundary-fixing isotopy. Hence, using the construction in Definition 3.4, we get the following result. In the statement, by trivial we mean that the vector field $X$ is of the form $\partial_{z}$ in a neighborhood of $\partial(\mathbb{D}\times[0,1])$, with $z\in[0,1]$.
Corollary 4.4.
For any $\epsilon\in\{\pm 1\}$, there exists a basic bordism
However, a computation step of a Turing machine is not only a shift, but also a read-write operation. We can represent this operation by gluing $W_{\epsilon}$ to suitable parts of disks. To be precise, suppose that we want to execute a transition $(\alpha,\beta,\epsilon)$, i.e., when the symbol $\alpha\in\mathcal{A}$ of the alphabet is read, the header substitutes it by $\beta\in\mathcal{A}$ and then shifts the tape in the $\epsilon=\pm 1$ direction.
For any symbol $s\in\mathcal{A}$, let us consider disjoint diffeomorphic open sets $\mathbb{D}_{0,s}$ of the disk $\mathbb{D}$ containing the vertical strip of all the points of $\Lambda$ of sequences $t$ with $t_{0}=s$, as in Figure 8. Using the shifting diffeomorphism $\varphi_{+}:\mathbb{D}\to\mathbb{D}$ of Lemma 4.3, we also set $\mathbb{D}_{n,s}=\varphi_{-}^{n}(\mathbb{D}_{0,s})$ for any $n\in\mathbb{Z}$. In this manner, the open set $\mathbb{D}_{n,s}$ contains the sequences $t\in\Lambda$ with $t_{n}=s$, see Figure 8.
Figure 8. Open sets $\mathbb{D}_{n,s}$.
Now, we are going to paste $W_{\epsilon}$ to an incoming disk $\mathbb{D}$ and to an outgoing disk $\mathbb{D}^{\prime}$ to represent the read-write-shift operation $(\alpha,\beta,\epsilon)$. For this purpose, let us denote the incoming boundary of $W_{\epsilon}$ by $\mathbb{D}(W_{\epsilon})$ and the outgoing boundary by $\mathbb{D}^{\prime}(W_{\epsilon})$. The partition defined above determines open sets $\mathbb{D}(W_{\epsilon})_{0,\beta}$ and $\mathbb{D}^{\prime}(W_{\epsilon})_{-\epsilon,\beta}$ of $\mathbb{D}(W_{\epsilon})$ and $\mathbb{D}^{\prime}(W_{\epsilon})$, respectively. Then, we glue a small cylinder $\mathbb{D}(W_{\epsilon})_{0,\beta}\times[0,1]$ to $\mathbb{D}(W_{\epsilon})_{0,\beta}$ and analogously a cylinder $\mathbb{D}^{\prime}(W_{\epsilon})_{-\epsilon,\beta}\times[0,1]$ to $\mathbb{D}^{\prime}(W_{\epsilon})_{-\epsilon,\beta}$. The resulting space will be denoted $\widehat{W}_{\epsilon}$ and is depicted in Figure 9.
Figure 9. Gluing of the bordism $W_{\epsilon}$ with two small cylinders.
Now, the resulting space $\widehat{W}_{\epsilon}$ can be glued to the incoming disk $\mathbb{D}$ and to an outgoing disk $\mathbb{D}^{\prime}$ to represent the read-write-shift operation $(\alpha,\beta,\epsilon)$. This is done by gluing the open set $\mathbb{D}_{0,\alpha}$ of $\mathbb{D}$ to the incoming boundary of $\widehat{W}_{\epsilon}$, which is $\mathbb{D}(W_{\epsilon})_{0,\beta}$ and coincides with $\mathbb{D}_{0,\alpha}$ up to traslation. Analogously, we glue the open set $\mathbb{D}^{\prime}(W_{\epsilon})_{-\epsilon,\beta}$ of the outgoing boundary of $\widetilde{W}_{\epsilon}$ to the open set $\mathbb{D}_{-\epsilon,\beta}^{\prime}$. Since the vector field is trivial in a boundary of the gluing loci, it naturally extends to a smooth vector field on the resulting space. The resulting bordism is shown in Figure 10.
Figure 10. Gluing of the bordism $W_{\epsilon}$ with the disks $\mathbb{D}$ and $\mathbb{D}^{\prime}$.
The flow of the resulting CW-complex performs exactly the read-write-shift operation $(\alpha,\beta,\epsilon)$. Indeed, since only the open set $\mathbb{D}_{0,\alpha}\subseteq\mathbb{D}$ is glued, the operation only affects to those tapes $t$ with $t_{0}=\alpha$. This open set $\mathbb{D}_{0,\alpha}$ is glued to $\mathbb{D}(W_{\epsilon})_{0,\beta}$, having the effect of substituting the symbol $\alpha$ at position 0 by the symbol $\beta$. In other words, under this gluing, a tape $t\in\mathbb{D}_{0,\alpha}\cap\Lambda$, i.e., with $t_{0}=\alpha$, corresponds to another tape $t^{\prime}\in\mathbb{D}(W_{\epsilon})_{0,\beta}\cap\Lambda$ with $t^{\prime}_{n}=t_{n}$ for $n\neq 0$ and $t^{\prime}_{0}=\beta$. Now, given such a tape $t^{\prime}\in\mathbb{D}(W_{\epsilon})_{0,\beta}\cap\Lambda$, the effect of the flow of $W_{\epsilon}$ is to send it to the tape $s_{\epsilon}(t^{\prime})\in\mathbb{D}^{\prime}(W_{\epsilon})\cap\Lambda$. Concretely, since $t^{\prime}_{0}=\beta$, then $s_{\epsilon}(t^{\prime})_{-\epsilon}=\beta$ and thus $s_{\epsilon}(t^{\prime})\in\mathbb{D}^{\prime}(W_{\epsilon})_{-\epsilon,\beta}$. This is exactly the open set that is glued identically with the final disk $\mathbb{D}^{\prime}$. Furthermore, the flow obtained in the resulting CW-complex is volume preserving, because the flows of each one of the basic bordisms that are glued are volume preserving as well.
Proposition 4.5.
For any $\epsilon\in\{\pm 1\}$, symbols $\alpha,\beta\in\mathcal{A}$, and disks $\mathbb{D}$ and $\mathbb{D}^{\prime}$, there exists a dynamical bordism endowed with a volume-preserving vector field that is trivial in a neighbourhood of the boundary such that the flow performs the read-write-shift operation $(\alpha,\beta,\epsilon)$ from the disk $\mathbb{D}$ to the disk $\mathbb{D}^{\prime}$.
4.3. Computational thickening
Consider a partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$. By [23], there exists a Turing machine $M=(Q,q_{0},Q_{\textrm{halt}},\delta)$ computing $f$. Furthermore, by Proposition 4.2, we can suppose that $M$ is reversible. In this section, we will prove that there exists a dynamical bordism $(W,X):\mathbb{D}\to\mathbb{D}$ such that $\psi_{(W,X)}:\Lambda\to\Lambda$ is the output function of $M$ (recall that the map $\psi_{(W,X)}$ was introduced in Section 3.3).
  • There is a single halting state node $q_{f}$. If there are more than one halting state nodes, then create a new halting state node and relabel all the former halting states as regular states. Also, create a new node with incoming edges of the form $(s,s,+)$ for all $s\in\mathcal{A}$, connecting all the formerly halting states with the new node, and outgoing vertices of the form (s,s,)($s,$-) connecting it with the new halting state, as in Figure 11.
    $q_{f_{1}}$$q_{f_{2}}$$e_{1}$$e_{2}$
    $q_{1}$$_{2}qq^{\prime}$$q_{f}$$e_{1}$$e_{2}$($s,s,+)($$s,s,+)($s,s,)($s,$-)
After these modifications, the proof will consist in ‘thickening’ the graph $\mathcal{G}_{M}$ to turn it into a bordism, with disks representing the states and tubes capturing the dynamics of the arrows. In these disks, we will encode the current state of the tape $t$ of the Turing machine $M$ as a point of $\Lambda\subseteq\mathbb{D}$, as explained in Section 3.2.
  1. (1)
    We place a disk for each vertex of $\mathcal{G}_{M}$. We denote the disk associated to the state $q\in Q$ by $\mathbb{D}(q)$.
  2. (2)
    Consider an edge in $\mathcal{G}_{M}$ between vertices $q$ and $q^{\prime}$ and with label $(\alpha,\beta,\epsilon)$, where $\alpha$ is the read symbol, $\beta$ is the written symbol, and $\epsilon=\pm 1$ is the shift. We glue the bordism $W_{\epsilon}$ to $\mathbb{D}(q)$ and $\mathbb{D}(q^{\prime})$ to represent the operation $(\alpha,\beta,\epsilon)$ as in Section 4.2, cf. Proposition 4.5.
This observation, together with the fact that the vector field is trivial around every boundary, implies that the vector fields on each of the bordisms glue together to define a smooth volume-preserving vector field $\widetilde{X}_{M}$ on $\widetilde{W}_{M}$. To construct the desired vector field on the homotopically equivalent bordism $W_{M}$, we use an auxiliary smooth manifold with boundary $\widehat{W}_{M}$, which is simply a small thickening of $\widetilde{W}_{M}$ where we round off the corners. It is enough to “smooth out” $\widetilde{X}_{M}$ only in a neighborhood of the disks, which allows us to define a volume preserving vector field $\widehat{X}_{M}$ on $\widehat{W}_{M}$ as the natural extension of $\widetilde{X}_{M}$ using that it is trivial around the boundary disks.
Next, we claim that the vector field $\widehat{X}_{M}$ can be extended to the whole bordism $W_{M}$ as a volume preserving vector field $X_{M}$. Indeed, since $\widehat{X}_{M}$ is volume preserving and it has zero flux across the boundary of the smooth manifold $\widehat{W}_{M}$ (by Gauss theorem), Poincaré’s lemma [31, Corollary 3.2.4] implies that there exists a smooth vector field $Y$ such that $\widehat{X}_{M}=\textup{curl}(Y)$.
[…]
Figure 12 depicts the smooth bordism $(W_{M},X_{M})$ obtained after the thickening of the original CW-complex $(\widetilde{W}_{M},\widetilde{X}_{M})$. In this figure, red disks represent outgoing boundaries of the basic bordisms, whereas blue disks are incoming boundaries. Observe that the red disks are disjoint, since only one transition occurs for every input, and similarly the blue disks are disjoint since the Turing machine is reversible. However, red disks are not disjoint from blue disks, but they are glued at different sides of the bigger black disks. Indeed, the computational dynamics arise precisely in these intersections between incoming and outgoing boundaries.
By construction, $W_{M}$ is a bordism with boundaries $\mathbb{D}(q_{0})$ and $\mathbb{D}(q_{f})$ corresponding to the disks attached to the (unique) starting and halting states of $M$. The flow of $X_{M}$ in $W_{M}$ satisfies the following properties.
First, given a state $q$, the points belonging to the set $\Lambda\cap\mathbb{D}(q)$ encode the tape state of the Turing machine $M$ when placed at state $q$ as a two-sided binary sequence. Furthermore, suppose that $x\in\Lambda\cap\mathbb{D}(q)$ represents a tape $t$. Let $\alpha=t_{0}$ be the initial symbol of the tape and $(q^{\prime},\beta,\epsilon)=\delta(q,\alpha)$ the resulting state $q^{\prime}$, symbol $\beta$ to write and shifting $\epsilon$ of the transition function of $M$ at this state. This means that there is an edge of the form $(\alpha,\beta,\epsilon)$ that joins $q$ and $q^{\prime}$ in the state machine of $M$. Then $x$ is glued to a bordism between $\mathbb{D}(q)$ and $\mathbb{D}(q^{\prime})$ that represents this transition and, as shown in Section 4.2, under this flow it reaches a unique point $x^{\prime}\in\Lambda\cap\mathbb{D}(q^{\prime})$ representing the resulting tape after executing the transition.
Now, let $x\in\Lambda\cap\mathbb{D}(q)$ be a point representing a tape $t$ that encodes a natural number $n\in\mathbb{N}$ in binary. Let $t^{0}=t,t^{1},t^{2},\ldots$ be the sequence of tapes obtained after successively executing $M$, which passes through states $q_{0},q_{1},q_{2},\ldots$ subsequently. By the previous observation, this uniquely corresponds to a sequence of points encoding the tapes $x_{0}=x,x_{1},x_{2},\ldots$, with $x_{i}\in\Lambda\cap\mathbb{D}(q_{i})$ for all $i$, such that the flow of $X_{M}$ connects them exactly in that order. Then we have two options:
  • If the Turing machine $M$ eventually stops, then $q_{N}=q_{f}$ for some $N$ large enough, and thus $x_{N}\in\Lambda\cap\mathbb{D}(q_{f})$ is in the outgoing boundary of $W_{M}$ and encodes the resulting tape $t^{N}$. By the construction of the Turing machine described in Section 4.1, the non-negative part of $t^{N}$ encodes the output $f(n)$ of the partial recursive function $f$ on $n$, whereas the negative part encodes the input $n$ itself.
$\psi_{(W_{M},X_{M})}:\mathbb{N}^{2}\dashrightarrow\mathbb{N}^{2}$
@@ 5. Categorical interpretation: Topological Kleene Field Theories
  • 22-Morphisms: Given two morphisms $f:\mathbb{N}^{\sqcup m}\dashrightarrow\mathbb{N}^{\sqcup m^{\prime}}$ and $g:\mathbb{N}^{\sqcup m}\dashrightarrow\mathbb{N}^{\sqcup m^{\prime}}$, a unique $2$-morphism $f\Rightarrow g$ between them exists if $f$ extends $g$, i.e., if $D_{g}\subseteq D_{f}$ and $f|_{D_{g}}=g$.
@@
In the same spirit, fixed $n\geq 2$, we can consider the category $\textup{Bord}_{n}^{\textup{dy}}$ of clean dynamical $n$-bordisms as the category given as follows:
@@
  • Morphism: A morphism $(M_{1},X_{1},\bm{\iota}_{1})\to(M_{2},\bm{\iota}_{2})$ between marked dynamical germs is an equivalence class of clean dynamical bordisms $(W,X)$ between them. Two such bordisms $(W,X)$ and $(W^{\prime},X^{\prime})$ are declared to be equivalent if there exists a boundary preserving diffeomorphism $F:W\to W^{\prime}$ such that $F_{*}(X)=X^{\prime}$.
@@
$Z:\mathcal{B}\longrightarrow\textup{PRF}.$
@@
is subjective or, in other words, any partial recursive function is representable through a dynamical bordism. Finally, the condition of being subjective on objects reduces to the fact that $\mathcal{B}$ contains at least one connected object.
Remark 5.7.
A way of constructing the category $\mathcal{B}$ of Definition 5.5 is by using a certain ‘core’ subcategory $\mathcal{B}_{0}$ of $\mathcal{B}$. This category $\mathcal{B}_{0}$ contains a single object $M=(M,X,\bm{\iota})$, with $M$ connected. It must also have the following properties:
  • For any partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, there exists an endobordism $(W,X):M\to M$ such that $Z_{0}(W,X)=f$.
  • Any endobordism $(W,X):M\to M$ of $\mathcal{B}_{0}$ satisfies that $Z_{0}(W,X)$ is a partial recursive function.
With this information at hand, $\mathcal{B}$ can be constructed as the monoidal closure of $\mathcal{B}_{0}$, i.e. the category made of disjoint unions of objects and bordisms of $\mathcal{B}_{0}$. Since $Z_{0}(M^{\sqcup n})=1^{\otimes n}=n$, we have that $Z=Z_{0}|_{\mathcal{B}}$ is subjective on objects and, since $Z_{0}|_{\mathcal{B}_{0}}$ takes values in PRF and PRF is monoidal, the functor $Z$ also takes values in PRF. Finally, by the very construction of $\mathcal{B}_{0}$, the functor $Z$ is full.
In this categorical language, Theorem 3.9 can be rephrased as follows.
@@
Topological Kleene Field Theories exist.
@@
By Theorem 3.9, given a partial recursive function, there exists a dynamical endobordism of the standard disk representing it. Now, consider the category $\mathcal{B}_{0}$ with $\mathbb{D}$ as the only object. For the morphisms, pick a Turing machine $M_{f}$ computing a partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$ for any $f$, and take as morphisms the compositional closure collection of the bordisms $(W_{M_{f}},X_{M_{f}})$ satisfying $Z_{0}(W_{M_{f}},X_{M_{f}})=f$. By construction, $\mathcal{B}_{0}$ satisfies the conditions of Remark 5.7 and thus defines a TKFT, proving the result. ∎
@@
We finish this section by pointing out that, to the best of our knowledge, it is not known whether the image of the functor $Z_{0}:\textup{Bord}_{n}^{\textup{dy}}\to\textup{PF}$ is always a partial recursive function. In other words, the clean condition on bordisms might force the function represented by a clean dynamical bordism to be computable. This would imply that the subcategory $\mathcal{B}$ of Definition 5.5 can be taken to be the full $\textup{Bord}_{n}^{\textup{dy}}$.
Theorem 5.9.
The reaching function of any clean dynamical bordism is a partial recursive functor. In other words, there exists a monoidal symmetric functor of $2$-categories
$Z:\textup{Bord}_{n}^{\textup{dy}}\to\textup{PRF},$
with values in the category of partial recursive functions.
@@ 6. Conclusions
The work presented in this paper establishes a new connection between dynamical systems and computability by representing computable functions through the flows of vector fields on bordisms. These ideas raise interesting questions to be explored in future research. Indeed, as a result of this work, we have proven that TKFTs represent a fully functional model of computation, on the same footing as Turing machines and partial recursive functions. However, the dynamical nature of the construction enables new perspective to address classical problems of computation through a novel glass.
Firstly, TKFTs also preclude that some functions might be speeded up using dynamical bordisms. It is well known that some beyond-Turing methods, such as Turing machines with advice, can be represented as area-preserving diffeomorphisms of the disk (cf. [10, Corollary 3.3]). In this sense, we expect that some “unconventional” diffeomorphisms of the disk may be used to speed-up the computation, the same way the advice accelerates the computation in super-Turing models. This emerging feature of clean dynamical bordisms would demonstrate that TKFT already exceeds the efficiency of Turing machines.
Directly related to this problem, the method developed in this paper links the complexity of a computable function with the topology of the bordism representing it. As is evident from the proof, a pair-of-pants, as a bordism between one disk and two disks, should be understood as an ‘if’ instruction in the computation, whereas a nontrivial handle represents a ‘loop’ construction. In fact, the method of proof creates a bordism that has as strong deformation retract the graph underlying a Turing machine computing the function, so the homology invariants of the bordisms, and in particular its first Betti number, are forced to agree with the ones of the graph. In this direction, a question arises: is it possible to (algorithmically) construct a clean dynamical bordism computing a partial recursive function $f$ with strictly smaller first Betti number than that of any Turing machine computing $f$? This is particularly interesting taking into account that TKFTs are equivalent to Turing machines, since using clean bordisms we will be able to compute exactly the same partial recursive functions, but potentially improving the complexity of the calculation.